QUESTION 4 Use the compound interest formulas A-P1.nt and A- Pelt to solve. Find the accumulated value of an investment of $1130 at 8% compounded annually for 14 years O $2395.60 $3073.17 $2305.20 $3319.03 Click Save and Submit to save and submit. Click Save All Answers to save all answers. hp ts 6 fe fg
Answer #1

A=P\left(1+\frac{r}{n}\right)^{nt}

here

P=1130

r=8% = 0.08

n=1 for annual compound

t=14 years

A=?

A=1130\left(1+\frac{0.08}{1}\right)^{1\cdot 14}

A=1130\left(1.08\right)^{14}

A=1130\cdot \:2.93719

{\color{Green} {\color{Red} A=3319.03}}

last option is correct